Calculate the force necessary to accelerate a 20,000-lbm rocket vertically upward at a rate of 100 ft/sec^2. Assume g 32.2 ft/sec 2
Answer #1

answer) we know the formula

F=ma=20000*100=2000000lbm-ft/s2

so necessary force=2000000/32.2=62112lbf.

so the answer is 62112lbf.